THE VECTOR SUM OF THE FORCES OF 10N and 6N CAN BE.
(a)2N
(b)8N
(c)18N
(d)20N

Grade:11

9 Answers

James Macmillan
11 years ago

ans (B) 8N

Beacause the maximum value of the magnitude of resultant vector of 10N and 6N ( 10 + 6 = 16N) can be 16N if the angle between the vector is 0 degree and the minimum value of magnitude of resultant vector can be 10 - 6 = 4N if the angle between the vector is 180 degree

Raunak Sudeep Taysheti
11 years ago
max=10+6 = 16N
min = 10-4 = 4 N
hence only option C is satisfying the conditions.
please apporve my answer by clicking on like.
Abhishekh kumar sharma
11 years ago

8N as net force lies between  lies between the range 4N and 16N

IT CAN BE EQUAL TO 2N FOR A SMALL ANGLE BETWEEN THEM

Akash Kumar Dutta
11 years ago

sum is given by root( A^2 + B^2 + 2.A.Bcosx)=(136 + 120cosx)
Hence ans is (b)8N ans.

tanvi khandelwal
10 years ago
the answer of this question is option(b) 8N because i have read this question in my resonance module.By- tanvi khandelwal
Rajwrita Nath
7 years ago
8
Tushar
5 years ago
The answer is 2 newtons I think this is the right answer because its not asking minimum or maximum so according to me the right answer is A
Pranav
5 years ago
From the rules of vector
A-B≤R≤A+B
 
Here, A = 10 and B= 6. So, 
10-6≤R≤10+6
4≤R≤16
 
Here, R is the resultant vector.So it should be below 16 and greater than 4. And from the option the favourable answer is 8 
 
So option (B) is correct!!!
 
Cheers!
Riya Pannase
5 years ago
Answer to this question is option b. 
   10+6=16 and 10-6=4. Therefore maximum is 16 and minimum is 4.  So answer will lie between 16 and 4.